Help please!
A survey of 100 college faculty shows that 67 own a car, 22 own a bicycle, and 11 own both. Find the probability that a faculty member chosen at random owns a car or a bicycle. (to two decimal places.)

Found 2 solutions by jorel1380, Maths68:
Answer by jorel1380(3719)

67+22-11=67+11=78
Probability=78/100=.78..

Answer by Maths68(1474)

P(own a car)=P(C)=67/100=0.67
P(own a bicycle)=P(B)=22/100=0.22
P(own car & bicycle) = P(C∩B) = 11/100=0.11
P (own car or bicycle) P(CUB)=?
P(CUB)=P(C)+P(B)- P(C∩B)
=(0.67+0.22)-0.11
=0.89-0.11
=0.78